# Parameter uncertainties for imperfect surrogate models in the low-noise regime

### v3

@@ III.3 High-dimensional models for atomic simulations
The results show that the envelope violation on test data varies between $2\%$ and $0.2\%$ for $N/P=0.06$ and $33$, respectively, in qualitative agreement with the results shown in figure 2. This demonstrates that worst-case model errors are extremely well captured by the POPS-hypercube ensemble. As reported in the left panel of figure 3, these excellent bounds do not trivially result from the resampled error distribution being overly pessimistic, as the distribution of errors induced by a uniform parameter resampling within the POPS-hypercube provides an excellent predictor of the actual error distribution with respect to the minimum-loss model. The MAE estimated from the resampled models exceeds the measured value by only 23%, while providing a very good description of the overall error distribution.
The information generated by this simple and efficient approach will be invaluable to estimate the uncertainties on a range of different physical quantities inferred from atomistic simulations.
IV Conclusion
@@
Synthetic test data
All synthetic data was generated with Python/JAX code, to be made available following peer review.
Atomic simulation data
Atomic training data was generated using an information-theoretic approach (Karabin and Perez, 2020; Montes de Oca Zapiain et al., 2022). 10,000 atomistic configurations containing between 2 and 32 tungsten atoms were generated and characterized using Density Functional Theory (Kohn and Sham, 1965), yielding 167,922 training data points in total (10,000 energies and 157,922 gradient components). Per-atom energies are expressed as a linear combination of $P=1596$ so-called bispectrum components within the ‘Quadratic’ Spectral Neighborhood Analysis Potential (SNAP) formalism (Wood and Thompson, 2018), which can be expressed as a linear surrogate model $V_{i}={\bm{\Theta}}\cdot{\bf F}_{i}(\{{\bf x}_{j}|r_{ij}<r_{c}\})$.
@@ Acknowledgements
This work was initiated during a visit at the Institute for Pure and Applied Mathematics at the University of California, Los Angeles (supported by National Science Foundation (NSF) grant DMS-1925919) and continued at the Institute for Mathematical and Statistical Innovation, University of Chicago (supported by the NSF grant DMS-1929348). Their hospitality is graciously acknowledged. TDS gratefully acknowledges support from ANR grants ANR-19-CE46-0006-1 and ANR-23-CE46-0006-1, IDRIS allocation A0120913455 and an Emergence@INP grant from the CNRS. D.P. was supported by the Laboratory Directed Research and Development program of Los Alamos National Laboratory under project number 20220063DR. Los Alamos National Laboratory is operated by Triad National Security, LLC, for the National Nuclear Security Administration of U.S. Department of Energy (Contract No. 89233218CNA000001).
Appendix A PAC-Bayes concentration inequalities for regression
PAC-Bayes inequalities were originally designed for classification problems where the loss function is bounded to $[0,1]$. The extension to regression requires treatment of unbounded losses, which is discussed in detail in (Germain et al., 2016).
Determining the value of the upper bound for $C_{0}[\pi_{0}]$ in the PAC-Bayes concentration inequality (21) requires certain assumptions on the distribution of the error between the empirical and true expected loss (Germain et al., 2016). Following (Germain et al., 2016), we assume that this error distribution is unbounded but sub-normal, i.e. all error fluctuations are bounded by a normal distribution of variance $s^{2}$, giving $C_{0}[\pi_{0}]<s^{2}/2$. We refer the reader to (Germain et al., 2016) and subsequent works (Shalaeva et al., 2020; Masegosa, 2020) for further discussion. For completeness, we recall here the bound used in the main text:
@@ Appendix B Model selection criteria for deterministic models
The Bayesian information criterion (BIC), used to discriminate between possible surrogate models, is derived by approximating (twice) the negative log evidence as $N/P\to\infty$, assuming a slowly varying prior $\pi_{0}({\bm{\Theta}})=\mathcal{O}(1)$. Using the above definitions we find
@@
We note that the value of the PAC-Bayes loss bound (21) satifies $BIC=2N\hat{\mathcal{L}}_{N}[\pi_{{\mathcal{L}},N}]-2C_{0}$, with $C_{0}$ model independent, as discussed in (Germain et al., 2016). In standard derivations ${\bm{\Sigma}}_{\mathcal{Y}}$ is set to the variational maximum likelihood solution ${\bm{\Sigma}}_{\mathcal{Y}}={\bm{\Sigma}}^{*}_{\mathcal{L}}$, yielding the familar expression $BIC=P\ln N+N\ln\|{\bm{\Sigma}}^{*}_{\mathcal{L}}\|+\ln\|{\boldsymbol{\mathcal{I}}}^{*}_{\mathcal{L}}\|$, where model independent terms are neglected. However, for deterministic regression problems, ${\bm{\Sigma}}_{\mathcal{Y}}$ is fixed. Neglecting model-independent terms, this gives an alternative BIC for deterministic models of
$BIC=N{\rm Tr}({\bm{\Sigma}}^{*}_{\mathcal{E}}{\bm{\Sigma}}^{-1}_{\mathcal{Y}})+P\ln N+\ln\|{\boldsymbol{\mathcal{I}}}^{*}_{\mathcal{L}}\|.$ (23)
@@ Appendix C Summary of Laplace’s method
Laplace’s method, also known as the steepest descents method, is a well-known identity allowing the evaluation of an integral of an exponentiated function multiplied by a large number. The method applies to a function $f(x)$ which is twice differentiable and has a unique maximum in some closed interval $[a,b]$ which may be the entire real line. This encompasses all applications in this paper; the extension to multidimensional functions is straightforward. Laplace’s method then reads
$\lim_{N\to\infty}\int_{a}^{b}\exp[Nf(x)]{\rm d}{x}=\sqrt{\frac{2\pi}{N}}\sqrt{\frac{{\rm d}^{2}f((x)}{{\rm d}x^{2}}}^{-1}\exp[N\max_{x\in[a,b]}f(x)].$ (24)
@@ Appendix D POPS-constrained ensemble ansatz
For any choice of weights $\mathcal{W}_{E}$, an approximate optimal set ${\bm{\Theta}}^{*}_{\mathbf{X}}\in\mathcal{T}^{*}_{E}$ can be determined in a variational setting at finite $N$, using (25) in the PAC-Bayes loss bound (21) for the loss ${\mathcal{L}}[\pi_{E}]=\langle{\rm w}_{E}({\mathbf{X}})L({\bm{\Theta}}_{\mathbf{X}})\rangle_{\mathcal{X}}$. As $P/N\to 0$ we show in the sub-section below this reduces to the set of POPS-constrained loss minimizers
@@
The POPS-constrained loss-minimization ensures the fits will be as close as possible under the Fisher metric ${\boldsymbol{\mathcal{I}}}^{*}_{\mathcal{L}}$, such that ${\bm{\Theta}}^{*}_{\mathbf{X}}\to{\bm{\Theta}}^{*}_{\mathcal{L}}$ for specified models. Under misspecification, the generalization error of (25) using $\mathcal{T}^{*}_{E}$ has the $P/N\to 0$ limit
${\mathcal{G}}_{0}[\pi_{E}]=-\left\langle\ln\left\langle{\rm w}_{E}({\mathbf{X}}^{\prime}){\rm m}_{\mathcal{E}}({\bm{\Theta}}^{*}_{{\mathbf{X}}^{\prime}},{\mathbf{X}})\right\rangle_{{\mathbf{X}}^{\prime}\in{\mathcal{X}}}\right\rangle_{{\mathbf{X}}\in{\mathcal{X}}}.$ (27)
In general we expect a uniform weighting ${\rm w}_{E}({\mathbf{X}})=1$ to be sub-optimal for generalization as the loss-minimizing fits ${\bm{\Theta}}^{*}_{\mathbf{X}}$ will be too close to ${\bm{\Theta}}^{*}_{\mathcal{L}}$.
@@
$\pi^{*}_{E}({\bm{\Theta}})\equiv\langle{\rm w}^{*}_{E}({\mathbf{X}})\delta({\bm{\Theta}}-{\bm{\Theta}}^{*}_{{\mathbf{X}}})\rangle_{\mathcal{X}}..$ (30)
Equation (30) is a variationally optimized ensemble ansatz for misspecified regression problems. The worst case of extreme misspecification, corresponding to non-intersecting POPS ${\mathcal{P}}_{\mathcal{M}}({\mathbf{X}}^{\prime})\cap{\mathcal{P}}_{\mathcal{M}}({\mathbf{X}})=\emptyset$, gives $\rho_{{\mathbf{X}}^{\prime}}^{*}({\mathbf{X}})=\delta_{{\mathbf{X}}{\mathbf{X}}^{\prime}}$ and thus uniform weights ${\rm w}^{*}_{E}({\mathbf{X}})=1$. Interestingly, it is simple to see we have uniform weighting at large $\epsilon$ as $\rho_{{\mathbf{X}}^{\prime}}^{*}({\mathbf{X}})$ is uniform. For real systems we find ${\rm w}^{*}_{E}({\mathbf{X}})$ is highly non-uniform and smallest for points near to ${\bm{\Theta}}^{*}_{\mathcal{L}}$, as expected. However, numerical test reveal moments of the predicted ensemble errors are still underestimated, as the ”loss like” bounds still bring all points too close to the minimum loss value, which motivates the hypercube approach described in the text. However, the max/min bounds of the ensemble, which are insensitive to our choice of ${\rm w}({\bf X})$, remain robust, as shown in figure 1.

### v5

@@ IV.3 Application to machine learning interatomic potentials
The properties of the ensemble are then evaluated on a complementary set of 10,000 hold-out testing points. Error analysis is carried out in the context of the POPS-hypercube ansatz. Even on the full dataset, the analysis requires only on the order of 20 minutes on 1 CPU core. It is therefore extremely lightweight in addition to being easily parallelizable, enabling its use for extremely large datasets. The results show that the envelope violation on test data varies between $2\%$ and $0.2\%$ for $N/P=0.06$ and $33$, respectively, in qualitative agreement with the results shown in figure 2. This demonstrates that worst-case model errors are extremely well captured by the POPS-hypercube ensemble. As reported in the left panel of figure 3, these excellent bounds do not trivially result from the resampled error distribution being overly pessimistic, as the distribution of errors induced by a uniform parameter resampling within the POPS-hypercube provides an excellent predictor of the actual error distribution with respect to the minimum-loss model. The MAE estimated from the resampled models exceeds the measured value by only 23%, while providing a very good description of the overall error distribution.
Our second test case fits a simple SNAP potential to ab initio data for the NbMoTaW high-entropy alloy from Li et al.Li et al. (2020). In this case, we use a simple form of multi-component descriptor that is linear in the number of species, and additionally choose a small set of descriptor functions, giving a total of $P=120$ adjustable parameters to fit $N=43,250$ force observations of this chemically complex system, with a further $N=121248$ force observations held out for testing. As can be seen in figure 4, the POPS-hypercube shows excellent performance in bounding prediction errors and in estimating the test error distribution, albeit with a slight underestimation in the tails of the error distribution.
To investigate the ability of our POPS approach to capture these large-error predictions, we also plot a common metric for uncertainty quantification schemes, the correlation between predicted and observed errors. Whilst such a correspondence is not expected to be exact, as the observed error is a stochastic quantity, there is clear benfit in the ability to asses, with confidence, whether a given prediction is expected to have small or large error. In the present setting, an appropriate comparison is the absolute error of the loss minimizing model against the predicted by the hypercube posterior prediction. A well perfoming model should a) bound the absolute error from above, i.e. not predict low error when the actual error is high and b) have a max-min range which increases for points with large error. As shown in the right of figure 4, we see these properties are indeed observed in the HEA example. An in-depth study of our POPS apporach in active learning schemes will be the subject of future work.
Figure 4: Fitting a SNAPWood and Thompson (2018) interatomic potential to a NbMoTaW high-entropy alloy (HEA) training setLi et al. (2020). Far Left: representative HEA configurations. Left: test error histogram at $P=120$, $N/P=376$ for the minimum loss model (black) and predictions from the POPS-hypercube $\pi^{*}_{\mathcal{H}}$ (green) ansatz. MAE: mean absolute error relative to the minimum loss solution. EV: envelope violation, points lying outside of the max/min bound. Right: probability of envelope violation for the $\pi^{*}_{\mathcal{H}}$ ansatz with $N/P$. Far right: correlation of actual vs predicted error in the test set. See main text.
IV.4 Application to molecular cheminformatics models
As a final application, we apply the POPS scheme to a recently introduced linear graphlet modelTynes et al. (2024) for predicting molecular properties. Whilst superficially similar to the interatomic potential application, the graphlet features are quite distinct as they are designed to produce explanatory models that detail the contribution of different local bond environment to the total atomic energy rather than providing a smoothly differentiable interatomic potential. The graphlet features, which are essentially discrete counts of the different possible subgraphs of the molecular graph, have a much higher sparsity than the atomic descriptors employed in the previous examples.
Here, we consider the formation energies of around 50,000 small organic molecules from the well-known QM9 dataset of small organic moleculesRuddigkeit et al. (2012); Ramakrishnan et al. (2014). Restricting feature extraction to graphlets of maximum order 3 or 4 gave a total of $P=247$ or $P=1239$ features, respectively. Once again, the formation energy model is linear in the graphlet features. As shown in figure 5, the POPS-hypercube again provides an excellent prediction and bounding of test errors, with the probability of envelope violation decreasing with increasing $N/P$. A detailed study of the use of our POPS approach for linear models in a wide range of applications will be the subject of a separate publication.
Figure 5: Fitting a linear graphlet modelTynes et al. (2024) to predict energies from the QM9 datasetRuddigkeit et al. (2012). Left: representative small organic molecules. Center: test error histogram at $P=247$, $N/P=42$ for the minimum loss model (black) and predictions from the POPS-hypercube $\pi^{*}_{\mathcal{H}}$ (green) ansatz. MAE: mean absolute error relative to the minimum loss solution. EV: envelope violation, points lying outside of the max/min bound. Right: Probability of envelope violation for the $\pi^{*}_{\mathcal{H}}$ ansatz with $N/P$.
V Comparison against existing approaches for uncertainty quantification and propagation in atomic simulation
As discussed in the previous section, uncertainty quantification schemes for machine learning interatomic potentials are actively sought after to assess the robustness of conclusions drawn from atomic simulationsWen and Tadmor (2020); Li et al. (2018); Gabriel et al. (2021); Wan et al. (2021); Best et al. (2024); Kellner and Ceriotti (2024); Thaler et al. (2023). Existing approaches for uncertainty quantification can be broadly categorized into two types, those that aim to directly estimate the error distribution of the surrogate model, and those (such as the POPS approach) that aim to capture model errors through the determination of parameter uncertainties.
Whilst direct error estimation is central to active learning workflowsLi et al. (2015); Podryabinkin and Shapeev (2017); Vandermause et al. (2020), these estimates can not typically be used to propagate uncertainty through the entire simulation engine to predict uncertainty on quantities of interest. This is because the result of any atomic simulation workflow is some quantity of interest (e.g. formation energies, elastic constants, or relaxed structures). Each quantity of interest is an (implicit) function of the parameters of the chosen interatomic potential, and will in general vary with the parameters in a non-trivial manner due to the strong correlations in the simulation routine. For example, computing defect formation energies require comparing two very similar, and thus highly correlated simulations. A direct estimate of energy error for each configuration will in general significantly overestimate the error in the formation energy, as the correlation between the two structures is not accounted for.
As a result, schemes which estimate uncertainty in model parameters are desirable for uncertainty propagation. The propagation can be achieved if we can predict how simulations are affected by variation in parameter, either through i) implicit differentiation methodsMaliyov et al. (2024) ii) reweighting of existing simulations using e.g. thermodynamic perturbation theoryImbalzano et al. (2021), or iii) direct resampling simulations by drawing samples from the parameter posterior distribution Best et al. (2024). A detailed study combining the first two of these approaches (implicit differentiation and thermodynamic perturbation) with the POPS procedure will be the subject of future work.
The POPS regression procedure presented above is a simple and efficient approach which returns uncertainties on the parameters of misspecified models. To the best of our knowledge, existing schemes for uncertainty quantification use as their objective function some form of loss, and thus cannot account for misspecification in parameter uncertainties of the model. In this sense, the POPS scheme is unique in its ability to robustly capture the parameter uncertainty of misspecified linear models.
The most directly-comparable scheme to POPS is ensembling, which typically train an ensemble of models on random subsets of the training data, a process known as baggingBreiman (1996). Various forms of ensemble methods are a popular approach to quantify uncertainty in neural network modelsAbe et al. (2022) and have been successfully applied to uncertainty quantification for neural network interatomic potentialsImbalzano et al. (2021); Kellner and Ceriotti (2024); Vita et al. (2024).
It is known that bagging and ensemble methods more generally are most effective for high-capacity models in presence of appreciable aleatoric uncertaintyBreiman (1996). In the low-noise regime, i.e. with weak aleatoric uncertainty, ensemble uncertainty estimates can be expected to significantly underestimate the true uncertainty, even when using high-capacity models such as neural networks. For low-capacity models, such as the underparametrized linear models considered here, bagging methods are not expected to produce useful uncertainty estimates. In most settings, loss-based uncertainty estimates require significant calibrationImbalzano et al. (2021); Kellner and Ceriotti (2024); Vita et al. (2024), or the use of closely-related methods such as conformal predictionBest et al. (2024), which similarly require some form of calibration. In addition, all loss-based schemes should give vanishing parameter uncertainty in the large data limit, as discussed in section I. Whilst bagging schemes for neural networks do provide finite parameter uncertainty, this is predominantly due to the multi-modal loss landscape, and has unclear relationship to the true uncertainty and generalization error of the model. Indeed, even the POPS-ensemble ansatz was found to underestimate moments of the error distribution, motivating the hypercube ansatz. An interesting direction for future work is to investigate how the POPS coverage criterion, equation (16), can be used to systematically construct improved ensemble methods in the low-noise regime.
We end this section with a comment on the wider validity of the POPS method for atomic simulation. Clearly, the requirement to be in the underparametrised limit $N\gg P$, can in practice lead to a large amount of training data when $P$ is large. In addition, this condition implicitly assumes that the model is not applied in a strongly extrapolative regime, which is formally true for any ensemble methodKellner and Ceriotti (2024). Whilst we have not found this to be an issue in practical applications, this theoretical requirement is also common to ensemble methodsKellner and Ceriotti (2024). In practical tests we do not find this to be problematic as the POPS scheme will typically give an overestimate of model uncertainty in regions of low training data density, which is a robust means to quantify extrapolation in high dimensionZeni et al. (2022). Nevertheless, we recommend using one of the many extrapolation grades / outlier measures for high-dimensional linear modelsPodryabinkin and Shapeev (2017); Lysogorskiy et al. (2021); Goryaeva et al. (2020); Zeni et al. (2022) to ensure that the POPS conditions are satisfied.
VI Conclusion
@@
Importantly, POPS encode model error as misspecification uncertainty on model parameters, which is essential for the propagation of uncertainty to simulation resultsMaliyov et al. (2024); Imbalzano et al. (2021). The application of POPS parameter uncertainties to multi-scale uncertainty propagation will be the subject of forthcoming work. In addition, a clear extension of the POPS approach is to non-linear models, for interatomic potentialsLysogorskiy et al. (2021); Vita et al. (2024) or more widely in scientific machine learning. As the loss function is non-convex, the central challenge will be to efficiently generate an optimal set of POPS-constrained loss minimizers, which can be achieved analytically in the linear case.
Data availability
A Python implementation of the POPS-constrained linear regression algorithm, following the Scikit-learn linear_model APIPedregosa et al. (2011), is available on GitHub at https://github.com/tomswinburne/POPS-Regression.git, where we also provide a Jupyter notebook to reproduce a selection of results from the paper. The library can also be installed via the pip package manager via pip install POPSRegression.
@@ Acknowledgements
This work was initiated during a visit at the Institute for Pure and Applied Mathematics at the University of California, Los Angeles (supported by National Science Foundation (NSF) grant DMS-1925919) and continued at the Institute for Mathematical and Statistical Innovation, University of Chicago (supported by the NSF grant DMS-1929348). Their hospitality is gratefully acknowledged. TDS was supported by ANR grants ANR-19-CE46-0006-1 and ANR-23-CE46-0006-1, IDRIS allocation A0120913455 and an Emergence@INP grant from the CNRS. D.P. was supported by the Laboratory Directed Research and Development program of Los Alamos National Laboratory under project number 20220063DR. Los Alamos National Laboratory is operated by Triad National Security, LLC, for the National Nuclear Security Administration of U.S. Department of Energy (Contract No. 89233218CNA000001).
Appendix A Glossary of terms
Table 1 contains a glossary of the key algebraic symbols used in the paper.
Symbol Description
${\mathcal{E}}$ Simulation engine
${\mathcal{M}}$ Surrogate model
${\mathcal{X}}$ Input space
${\mathcal{Y}}$ Output space
${\mathbf{X}}$ Input vector
${\mathbf{Y}}$ Output vector
$X$ Input vector dimension
$Y$ Output vector dimension
${\bm{\Theta}}$ Model parameters
${\mathcal{P}}$ Parameter space
$N$ Number of training points
$P$ Number of model parameters
$\rho_{\mathcal{E}}({\mathbf{Y}}|{\mathbf{X}})$ Output distribution of simulation engine
$\rho_{\mathcal{M}}({\mathbf{Y}}|{\mathbf{X}},{\bm{\Theta}})$ Output distribution of surrogate model for given parameter choice
$\pi({\bm{\Theta}})$ Posterior parameter distribution
$\rho_{\mathcal{M}}({\mathbf{Y}}|{\mathbf{X}})$ Model output distribution $\rho_{\mathcal{M}}({\mathbf{Y}}|{\mathbf{X}},{\bm{\Theta}})$ averaged over $\pi({\bm{\Theta}})$
${\mathcal{G}}[\pi]$ Generalization error of $\rho_{\mathcal{M}}({\mathbf{Y}}|{\mathbf{X}})$ to $\rho_{\mathcal{E}}({\mathbf{Y}}|{\mathbf{X}})$ under $\pi({\bm{\Theta}})$
${\mathcal{L}}[\pi]$ Expected loss of $\rho_{\mathcal{M}}({\mathbf{Y}}|{\mathbf{X}})$ to $\rho_{\mathcal{E}}({\mathbf{Y}}|{\mathbf{X}})$ under $\pi({\bm{\Theta}})$
$\hat{L}_{N}({\bm{\Theta}})$ Empirical loss of $\rho_{\mathcal{M}}({\mathbf{Y}}|{\mathbf{X}},{\bm{\Theta}})$ to $\rho_{\mathcal{E}}({\mathbf{Y}}|{\mathbf{X}})$ with $N$ training points
${\bm{\mathcal{I}}}^{*}_{\mathcal{L}}$ $X\times X$ Fisher information matrix from empirical loss
${\bm{\Sigma}}_{\mathcal{Y}}$ $Y\times Y$ Aleatoric covariance matrix (assumed vanishing, determinant $\mathcal{O}(\epsilon^{2Y})$)
$\epsilon$ Dimensionless measure of aleatoric uncertainty (assumed small)
${\bm{\Sigma}}^{*}_{\mathcal{L}}$ $Y\times Y$ Error covariance matrix from loss minimization
$\pi_{0}({\bm{\Theta}})$ Prior parameter distribution
$\pi_{{\mathcal{L}},N}({\bm{\Theta}})$ Posterior parameter distribution from Bayesian inference
$\pi^{*}_{E}({\bm{\Theta}})$ Posterior parameter distribution from POPS-constrained ensemble ansatz
$\pi^{*}_{H}({\bm{\Theta}})$ Posterior parameter distribution from POPS-constrained hypercube ansatz
Table 1: Glossary of key algebraic symbols used in the paper
@@ Appendix B Model selection criteria for deterministic models
The Bayesian information criterion (BIC), used to discriminate between surrogate models, is derived by approximating (twice) the negative log evidence as $N/P\to\infty$, assuming a slowly varying prior $\pi_{0}({\bm{\Theta}})=\mathcal{O}(1)$. We have
@@
We note that the value of the PAC-Bayes loss bound (8) satisfies $BIC=2N\hat{\mathcal{L}}_{N}[\pi_{{\mathcal{L}},N}]-2C_{0}$, with $C_{0}$ model independent, as discussed in Ref. Germain et al., 2016. In standard derivations, ${\bm{\Sigma}}_{\mathcal{Y}}$ is set to the variational maximum likelihood solution ${\bm{\Sigma}}_{\mathcal{Y}}={\bm{\Sigma}}^{*}_{\mathcal{L}}$, yielding the familar expression $BIC=P\ln N+N\ln\|{\bm{\Sigma}}^{*}_{\mathcal{L}}\|+\ln\|{\bm{\mathcal{I}}}^{*}_{\mathcal{L}}\|$, where model independent terms are neglected (principally, terms containing only $N$ and the aleatoric covariance $\Sigma_{\mathcal{Y}}$). However, for deterministic regression problems, ${\bm{\Sigma}}_{\mathcal{Y}}$ is fixed. Neglecting model-independent terms, this gives an alternative BIC for deterministic models of
$BIC=N{\rm Tr}({\bm{\Sigma}}^{*}_{\mathcal{L}}{\bm{\Sigma}}^{-1}_{\mathcal{Y}})+P\ln N+\ln\|{\bm{\mathcal{I}}}^{*}_{\mathcal{L}}\|.$ (23)
In physical terms, this corresponds to an estimation of a partition function based on an harmonic approximation of the Hamiltonian around its minimum.
@@ Appendix C Summary of Laplace’s method
Laplace’s method, also known as the steepest descents method, is a well-known identity allowing the evaluation of an integral of an exponentiated function multiplied by a large number. The method applies to a function $f(x)$ which is twice differentiable and has a unique maximum in some closed interval $[a,b]$ which may be the entire real line. This encompasses all applications in this paper; the extension to multidimensional functions is straightforward. The method approximates the integral using a second-order Taylor expansion of $f(x)$ around the maximum of the function, which reduces the problem to a Gaussian integral. Laplace’s method then reads
$\lim_{N\to\infty}\int_{a}^{b}\exp[Nf(x)]{\rm d}{x}=\sqrt{\frac{2\pi}{N}}\sqrt{\frac{{\rm d}^{2}f(x^{*})}{{\rm d}x^{2}}}^{-1}\exp[Nf(x^{*})],\quad x^{*}=\arg\max_{x\in[a,b]}f(x).$ (24)
For further information we refer the reader to e.g. (Wong, 2001).
@@ Appendix D POPS-constrained ensemble ansatz
For any choice of weights $\mathcal{W}_{E}$, an approximate optimal set ${\bm{\Theta}}^{*}_{\mathbf{X}}\in\mathcal{T}^{*}_{E}$ can be determined in a variational setting at finite $N$, using (25) in the PAC-Bayes loss bound (8) for the loss ${\mathcal{L}}[\pi_{E}]=\langle{\rm w}_{E}({\mathbf{X}})L({\bm{\Theta}}_{\mathbf{X}})\rangle_{\mathcal{X}}$. As $P/N\to 0$, we show in the sub-section below that this reduces to the set of POPS-constrained loss minimizers
@@
The POPS-constrained loss-minimization ensures the fits will be as close as possible under the Fisher metric ${\bm{\mathcal{I}}}^{*}_{\mathcal{L}}$, such that ${\bm{\Theta}}^{*}_{\mathbf{X}}\to{\bm{\Theta}}^{*}_{\mathcal{L}}$ for specified models. Under misspecification, the generalization error of (25) using $\mathcal{T}^{*}_{E}$ has the $P/N\to 0$ limit
${\mathcal{G}}_{0}[\pi_{E}]=-\left\langle\ln\left\langle{\rm w}_{E}({\mathbf{X}}^{\prime}){\rm m}_{\mathcal{E}}({\bm{\Theta}}^{*}_{{\mathbf{X}}^{\prime}},{\mathbf{X}})\right\rangle_{{\mathbf{X}}^{\prime}\in{\mathcal{X}}}\right\rangle_{{\mathbf{X}}\in{\mathcal{X}}},$ (27)
where ${\rm m}_{\mathcal{E}}({\bm{\Theta}},{\mathbf{X}})$ is the mass function introduced in (14). In general we expect a uniform weighting ${\rm w}_{E}({\mathbf{X}})=1$ to be sub-optimal for generalization as the loss-minimizing fits ${\bm{\Theta}}^{*}_{\mathbf{X}}$ will be too close to ${\bm{\Theta}}^{*}_{\mathcal{L}}$.
@@
$\pi^{*}_{E}({\bm{\Theta}})\equiv\langle{\rm w}^{*}_{E}({\mathbf{X}})\delta({\bm{\Theta}}-{\bm{\Theta}}^{*}_{{\mathbf{X}}})\rangle_{\mathcal{X}}.$ (30)
Equation (30) is a variationally optimized ensemble ansatz for misspecified regression problems. The worst case of extreme misspecification, corresponding to non-intersecting POPS ${\mathcal{P}}_{\mathcal{M}}({\mathbf{X}}^{\prime})\cap{\mathcal{P}}_{\mathcal{M}}({\mathbf{X}})=\emptyset$, gives $\rho_{{\mathbf{X}}^{\prime}}^{*}({\mathbf{X}})=\delta_{{\mathbf{X}}{\mathbf{X}}^{\prime}}$ and thus uniform weights ${\rm w}^{*}_{E}({\mathbf{X}})=1$. Interestingly, it is simple to show that optimal weights are also uniform in the large $\epsilon$ limit, as $\rho_{{\mathbf{X}}^{\prime}}^{*}({\mathbf{X}})$ is then uniform. For real systems we find ${\rm w}^{*}_{E}({\mathbf{X}})$ is highly non-uniform and smallest for points near to ${\bm{\Theta}}^{*}_{\mathcal{L}}$, as expected. However, while this approaches improves the description of the parameter uncertainty compared to the uniform weight variant described in the main text, numerical tests reveal that moments of the predicted ensemble errors are still underestimated, as the ”loss like” bounds still bring all points too close to the minimum loss value. This observation (as well as the reduced evaluation cost) motivates the hypercube approach described in the text. However, the max/min bounds of the ensemble, which are insensitive to our choice of ${\rm w}({\bf X})$, remain robust, as shown in figure 1.